# Regularity of Einstein $5$-manifolds via $4$-dimensional gap theorems

### v1

@@ Abstract.
  • tangent cones are unique of the form $\mathbb{R}\times\mathbb{R}^{4}/\Gamma$ on the top stratum, hence outside a countable set of points,
@@
The proofs rely on a new result of independent interest: all spherical and hyperbolic $4$-orbifolds are isolated among Einstein $4$-orbifolds in the Gromov-Hausdorff sense. This yields various gap theorems for Einstein $4$-orbifolds, which do not extend to higher dimensions.
@@ 1. Introduction
In the present article, we refine our understanding of limits in dimension $4$ and deduce several gap theorems for spherical and hyperbolic orbifolds, only valid in this dimension. This leads to substantial applications to the regularity of limits of Einstein $5$-manifolds, towards an orbifold regularity theorem away from a codimension $5$ singular set.
@@ 1.1.1. Noncollapsed degeneration of Einstein $4$-manifolds
It has long been an open question, explicitly posed in [And94, And10], whether Einstein orbifolds can be approximated or resolved by sequences of smooth Einstein metrics. Such approximations have been obtained by gluing-perturbation, see e.g. [Biq13, Don12, LS94]. On the other hand, in [Ozu21b], it was proven that spherical and hyperbolic orbifolds with the simplest singularities, $\mathbb{R}^{4}/\pm$, could never be limits of smooth Einstein $4$-manifolds in the Gromov-Hausdorff sense. In the present article, we show that all hyperbolic and spherical orbifolds are $GH$-isolated, even within the class of Einstein $4$-orbifolds. This is not a mere extension of [Ozu21b] and requires completely new ideas in order to deal with the deeply technical case of multiscale trees of singularities and singular metrics. The resulting isolation theorem implies a series of new surprising quantitative gap phenomena for Einstein $4$-orbifolds. These results are genuinely $4$-dimensional: they have no analogue in higher dimensions $n\geqslant 5$, even with isolated singularities.
@@ 1.1.2. Regularity of noncollapsed Einstein $n$-manifolds for $n\geqslant 5$
The higher-dimensional Einstein manifolds have been extensively studied over the last four decades. The modern compactness theory originates with the Cheeger–Gromov program of pointed $C^{\infty}$ convergence under curvature control. A foundational leap was the structure theory for spaces with lower Ricci curvature bounds [CC96, Col97, CC97, CC00a, CC00b], which introduced the stratification of the singular set and established the regularity of the limit space.
@@
There is a substantial literature on this uniqueness problem of geometric PDEs. For minimal surface, the uniqueness was established assuming the integrability and smoothness of the cross-sections in [AA81], and the integrability assumption was later removed in the seminal work of [Sim81]. For more general nonsmooth cross-sections, the uniqueness of cones $\mathds{R}\times C(\mathbb{S}^{p}\times\mathbb{S}^{q})$ for $p+q\geq 7$ was proved in [Sim94]. Only very recently was the uniqueness of the cylindrical Simons cone, $\mathds{R}\times C(\mathbb{S}^{3}\times\mathbb{S}^{3})$ established in [Sze20a]. The case $\mathds{R}\times C(\mathbb{S}^{2}\times\mathbb{S}^{4})$ was subsequently addressed in [FTW25]. See also the uniqueness for two-dimensional area-minimizing integral currents [Whi83]. Analogous uniqueness problems arise in other geometric PDEs, for example, mean curvature flow [Sch14, CIM15, CM15, CS21, CM23, LZ24], and Ricci flow [SW15, CMZ24, LW24, CM25, FL25a] and the references therein.
For noncollapsed Ricci-flat manifolds, Cheeger-Tian [CT94] proved uniqueness under the assumption that the cross-sections of all tangent cones are smooth and integrable; later Colding-Minicozzi [CM14] established uniqueness assuming only that one tangent cone is smooth. See also the Kähler-Einstein setting [DS17] and related study [KSz24, YZ25]. When the cone has a nonsmooth cross-section, for instance a cone that splits off a line $\mathbb{R}\times C(Z)$, much less is known.
As our first result in higher dimensions, we show that in dimension 5, if one tangent cone splits off a line, then it is the unique tangent cone. To the best of authors’ knowledge, this is the first uniqueness result for a given general Ricci-flat cone with a nonsmooth cross-section. It yields a clearer understanding of the asymptotic geometry and provides curvature decay estimates for Ricci-flat $5$-manifolds.
Regularity of limit spaces and singular sets. Through a series of works [And89, BKN89, Tia90, CN15], it was shown that limits of $4$-manifolds with bounded Ricci curvature and uniform lower volume bounds are Riemannian orbifolds with locally finitely many singularities. A natural question is whether such structural results extend to higher dimensions. Key recent milestones include the metric geometry structure of the regular set, the resolution of the codimension-four conjecture and the finiteness of $\mathcal{H}^{n-4}$-measure of the singular set [CoN12, CN15, CJN21, JN21]. Moreover, the rectifiability of the singular set was established in [CJN21, JN21]: up to an $\mathcal{H}^{n-4}$-measure zero subset, $\mathcal{S}$ is contained in a countable union of bi-Lipschitz images of subsets of $\mathbb{R}^{n-4}$.
@@
    Refined rectifiability. We prove that the entire singular set is contained in a countable union of points and bi-Lipschitz images of subsets of $\mathbb{R}$, without discarding any $\mathcal{H}^{1}$-measure zero subset. This refinement relies on the uniqueness of local tangent cones: while [CJN21, JN21] established uniqueness $\mathcal{H}^{n-4}$-almost everywhere, in dimension $5$ we prove uniqueness on the entire top stratum, hence away from a countable collection of points.
@@
    Orbifold regularity. We extend the orbifold regularity to dimension $5$ and formulate a codimension-five orbifold conjecture for Einstein limits. In particular, the singular curves, away from a nowhere-dense subset $\mathcal{S}_{\mathrm{orb}}$, are geodesic and possess a real-analytic orbifold structures with bounded curvature. By classifying all geodesics in the limit space $(X,d)$, we also obtain more metric geometry properties of the limit space, generalizing [CoN12]. Namely, the orbifold-regular set $\mathcal{R}_{\mathrm{orb}}:=X\setminus\mathcal{S}_{\mathrm{orb}}$ is an open convex orbifold, and the orbifold-singular set $\mathcal{S}_{\mathrm{orb}}$ is the set of extreme points of the limit space. Since Einstein metrics are real-analytic, we further conjecture – by unique continuation – that $\mathcal{S}_{\mathrm{orb}}$ is countable. This will be discussed in detail in Section 1.3.
Many of the results above require only bounded Ricci curvature, and less central statements hold in all dimensions $n\geqslant 5$. See [Col96a, Col96b, CT06, Col12, CN13] for more developments in this direction.
@@ 1.2.1. Gap theorems for Einstein $4$-manifolds
There are Ricci-flat $4$-manifolds approaching flat orbifolds [LS94] on $K3$, but not on any topology [BK17, Ozu21a].
@@
    (Almost Lévy–Gromov theorem) there exists some $t\in(17^{-17},1-17^{-17})$ with $\big|I_{(M,g)}(t)-I_{\mathbb{S}^{4}}(t)\big|\leqslant\varepsilon_{0}$, where $I$ denotes the isoperimetric profile $I_{(M,g)}(t):=\inf\big\{\mathrm{Per}(\Omega):\Omega\subset M,{\rm Vol}_{g}(\Omega)=t\,{\rm Vol}_{g}(M)\big\},$
@@
Each gap theorem above strengthens well-known rigidity theorems for spherical suspensions. For (1) use the maximal-diameter theorem of Cheng [Che75], for (2) use the Lichnerowicz–Obata theorem [Lic58, Oba62], for (4) use the Lévy–Gromov estimate [Gro81] (see a survey [Ant25]), for (5) use Bishop–Gromov based at the singular point [BC64, Gro81]. In the RCD setting, equality yields a spherical suspension; see for instance [Ket15, CaM17]. Note also that the $L^{1}$ smallness in (3) is implied by the $L^{p}$ smallness with any $p\geqslant 1$ and that gap theorems for smooth Einstein $n$-manifolds with small $L^{\frac{n}{2}}$ norm of their Weyl tensors are proven in [Sin92, She90, IS02].
@@
Generalizations of Corollary 1.3 fail in dimension $5$ and higher: in [Böh98], Böhm constructed a sequence of Einstein $5$-metrics $(S^{5},g_{i})$ with ${\rm Ric}(g_{i})=4g_{i}$ and volume uniformly bounded below such that $(S^{5},g_{i})_{i\in\mathbb{N}}$ converges in the Gromov-Hausdorff sense to a spherical suspension where the diameter and the first eigenvalue of the $(S^{5},g_{i})$ are arbitrarily close to $\pi$ and $5$ respectively, and supposedly, since the singularities forming along the sequence are conical, we should have: for $1\leqslant p<\frac{5}{2}$, $\|W\|_{L^{p}}\to 0$.
@@
The constants in our gap theorems of Corollary 1.3 are not explicit since they are proven by contradiction. Gap theorems with explicit constant, such as [Gur00] are exceptionally rare. See related discussion of gap theorems in [HM25].
@@ 1.2.3. Singular set of noncollapsed limits of Einstein $5$-manifolds, and orbifold structure
where $\mathcal{S}_{\Gamma}^{1}$ is the set of points whose unique tangent cone is $\mathds{R}\times\mathds{R}^{4}/\Gamma$, $\Gamma\subset O(4)$ and $\mathcal{S}^{0}$ is a set of countably many points where any tangent cone never splits.
Our first result refines the classical rectifiability statement of [CJN21, JN21], eliminating the need to discard an $\mathcal{H}^{1}$-measure zero subset:
@@
Let $(M_{i},g_{i},p_{i})_{i\in\mathbb{N}}$ be a sequence of pointed $5$-manifolds with $|{\rm Ric}|<4$ and ${\rm Vol}(B_{1}(p_{i}))>v>0$. Let $(X,d,p)$ be a Gromov-Hausdorff limit of the sequence. Then the singular set $\mathcal{S}$ satisfies: There exist countably many points $q_{j}$ and measurable subsets $Z_{i}\subset\mathcal{S}^{1}_{\Gamma_{i}}$ with bi-Lipschitz maps $\phi_{i}:Z_{i}\to\mathds{R}$ such that
@@
Theorem 1.18 characterizes any curve of singularities up to a nowhere dense subset: they are open geodesic segments. We now turn to the converse and completely characterize neighborhoods of open geodesic segments of our limit spaces.
@@
Let $\sigma:[0,l]\to X$ be a geodesic in a noncollapsed Einstein $5$-dimensional limit space $(X,d)$ with $|{\rm Ric}_{M_{i}}|\leqslant 4$. Then either $\sigma\big((0,l)\big)$ is regular, or the entire $\sigma\big((0,l)\big)\subset\mathcal{S}^{1}_{\Gamma}$ for some $\Gamma\subset O(4)$ and the interior of the curve is removable in the sense of Theorem 1.18.
@@
Finally we discuss the orbifold regularity of the limit space $X$. Theorem 1.18 says the “interior” of each $\mathcal{S}^{1}_{\Gamma}$ is a countable union of geodesic segments along which the metric extends real-analytically. Then it is natural to introduce the orbifold-regular-singular decomposition as follows:
@@
1.2.4. Singular set of noncollapsed limits of Einstein $n$-manifolds
In higher dimensions our information about the singular locus is more limited: we can distinguish the (n4)(n-4)-stratum from the (n5)(n-5)-stratum, and this separation already forces a uniqueness phenomenon.

### v2

@@ Abstract.
  • tangent cones are unique of the form $\mathbb{R}\times\mathbb{R}^{4}/\Gamma$ on the top stratum, hence outside a countable set of points; this follows from a new isolation result for cones of the form $\mathbb{R}\times\mathbb{R}^{4}/\Gamma$ among all tangent cones,
@@
The proofs rely on a new result of independent interest: all spherical and hyperbolic $4$-orbifolds are isolated among Einstein $4$-orbifolds in the Gromov-Hausdorff sense. This yields various gap theorems for Einstein $4$-orbifolds, which do not extend to higher dimensions. The proofs of these gap theorems require a careful analysis of singular metrics and families of metrics that degenerate.
@@ 1. Introduction
In the present article, we refine our understanding of limits in dimension $4$ and deduce several gap theorems for spherical and hyperbolic orbifolds, only valid in this dimension. This leads to substantial applications to the regularity of limits of Einstein $5$-manifolds, towards an orbifold regularity theorem away from a conjectually codimension $5$ singular set.
@@ 1.1.1. Noncollapsed degeneration of Einstein $4$-manifolds
It has long been an open question, explicitly posed in [And94, And10], whether Einstein orbifolds can be approximated or resolved by sequences of smooth Einstein metrics. Such approximations have been obtained by gluing-perturbation, see e.g. [Biq13, Don12, LS94]. On the other hand, in [Ozu21b], it was proven that spherical and hyperbolic orbifolds with the simplest singularities, $\mathbb{R}^{4}/\pm$, could never be limits of smooth Einstein $4$-manifolds in the Gromov-Hausdorff sense. In the present article, we show that all hyperbolic and spherical orbifolds are $GH$-isolated, even within the class of Einstein $4$-orbifolds. This is not a mere extension of [Ozu21b] and requires completely new ideas in order to deal with the deeply technical case of multiscale trees of singularities and singular metrics. The resulting isolation theorem implies a series of new surprising quantitative gap phenomena for Einstein $4$-orbifolds. These results are genuinely $4$-dimensional: they have no analogue in higher dimensions $n\geqslant 5$, even with isolated singularities. This is the first instance of gap theorems that deals with orbifold degenerations, while previous ones such as [HM25] relied on topological or geometric assumptions that prevented said degenerations.
@@ 1.1.2. Regularity of noncollapsed Einstein $n$-manifolds for $n\geqslant 5$
The higher-dimensional Einstein manifolds have been extensively studied over the last four decades. The modern compactness theory originates with the Cheeger-Gromov program of pointed $C^{\infty}$ convergence under curvature control. A foundational leap was the structure theory for spaces with lower Ricci curvature bounds [CC96, Col97, CC97, CC00a, CC00b], which introduced the stratification of the singular set and established the regularity of the limit space.
@@
Such uniqueness for noncollapsed Ricci-flat manifolds was first established by Cheeger-Tian [CT94] under the assumption that the cross-sections of all tangent cones are smooth and integrable. Two decades later, Colding-Minicozzi [CM14] proved uniqueness assuming only that a single tangent cone is smooth, using the Łojasiewicz-Simon inequality. See also the related developments [KSz24, YZ25]. By contrast, when the cone has a nonsmooth cross-section – most notably for cones that split off a line $\mathbb{R}\times C(Z)$ – the theory remains far less developed. In the Kähler-Einstein setting, Donaldson-Sun [DS17] established uniqueness of tangent cones by means of algebro-geometric methods.
As our first result in higher dimensions, we prove that in dimension five, if one tangent cone splits off a line, then it is the unique tangent cone. Our argument does not rely on the Łojasiewicz-Simon inequality as in [CM14], nor on algebro-geometric techniques as in [DS17]. To the best of the authors’ knowledge, this is the first uniqueness result for a general Ricci-flat cone with a nonsmooth cross-section. As a consequence, we obtain a clearer description of the asymptotic geometry and new curvature decay estimates for Ricci-flat $5$-manifolds.
Analogous uniqueness questions arise in other geometric PDEs, including minimal surface theory [AA81, Sim83, Whi83, Sim94, Sze20b, CL23, FTW25], mean curvature flow [Sch14, CIM15, CM15, CS21, CM23, LZ24, BL25], and Ricci flow [SW15, CMZ24, LW24, CM25, FL25a], where uniqueness has been obtained under various smoothness, integrability, and structural assumptions on the cross-sections. We refer to the literature for further developments.
Regularity of limit spaces and singular sets. Through a series of works [And89, BKN89, Tia90, CN15], it was shown that limits of Einstein $4$-manifolds with bounded Ricci curvature and uniform lower volume bounds are Riemannian orbifolds with locally finitely many singularities. A natural question is whether such structural results extend to higher dimensions. Key recent milestones include the convexity of the regular set [CoN12], the resolution of the codimension-four conjecture [CN15] and the rectifiability of the singular set [CJN21, JN21]. Here, rectifiability means that, up to an $\mathcal{H}^{n-4}$-measure zero subset, $\mathcal{S}$ is contained in a countable union of bi-Lipschitz images of subsets of $\mathbb{R}^{n-4}$.
@@
    Refined rectifiability and convexity. We prove the uniqueness of the tangent cone on the entire top stratum: if one tangent cone is $\mathds{R}\times\mathds{R}^{4}/\Gamma$, then it is the unique tangent cone. The uniqueness yields a refined stratification of the singular set $\mathcal{S}$ as a finite disjoint union of sets $\mathcal{S}^{0}$ and $\mathcal{S}^{1}_{\Gamma}$ for finitely many $\Gamma\subset O(4)$, where $\mathcal{S}^{0}$ consists of countably many points whose tangent cones do not split, and $\mathcal{S}^{1}_{\Gamma}$ consists of points whose tangent cone is $\mathbb{R}\times\mathbb{R}^{4}/\Gamma$. We show that each $\mathcal{S}^{1}_{\Gamma}$ is locally contained in a bi-Lipschitz image of a subset of $\mathbb{R}$, thereby establishing a bi-Lipschitz manifold structure for the singular set without discarding any $\mathcal{H}^{1}$-measure zero subset. Moreover, we prove that each connected component of $\mathcal{S}^{1}_{\Gamma}$ is either a single point or a geodesic, and hence locally convex.
@@
    Orbifold regularity. We extend the orbifold regularity to dimension $5$ and formulate a codimension-five orbifold conjecture for Einstein limits $X$. In particular, the singular curves in $\mathcal{S}^{1}_{\Gamma}$ are geodesic and possess a real-analytic orbifold structures with bounded curvature. Together with the geodesics in the regular set, these exhaust all geodesics in the limit space. This leads us to introduce the orbifold-regular-singular decomposition $X=\mathcal{R}_{\mathrm{orb}}\cup\mathcal{S}_{\mathrm{orb}}$. We prove that the orbifold-regular set $\mathcal{R}_{\mathrm{orb}}$ is an open convex 5-dimensional orbifold, and the orbifold-singular set $\mathcal{S}_{\mathrm{orb}}$ consists precisely of the extreme points of $X$, namely those points that do not lie on any geodesic joining two other points. We further show that $\mathcal{S}_{\mathrm{orb}}$ is contained in a nowhere dense subset of a countable collection of bi-Lipschitz curves. Since Einstein metrics are real-analytic, we further conjecture – by unique continuation – that $\mathcal{S}_{\mathrm{orb}}$ is countable. This will be discussed in detail in Section 1.3.
Many of the results above require only bounded Ricci curvature, and less central statements hold in all dimensions $n\geqslant 5$. See [Col96a, Col96b, CCT02, CT06, Col12, CN13, Bam17, Bam18, Bam20] for more developments in this direction.
@@ 1.2.1. Gap theorems for Einstein $4$-manifolds
There are Ricci-flat $4$-manifolds approaching flat orbifolds [LS94] on $K3$, but not on arbitrary topologies [BK17, Ozu21a].
@@
    (Almost Lévy-Gromov theorem) there exists some $t\in(17^{-17},1-17^{-17})$ with $\big|I_{(M,g)}(t)-I_{\mathbb{S}^{4}}(t)\big|\leqslant\varepsilon_{0}$, where $I$ denotes the isoperimetric profile $I_{(M,g)}(t):=\inf\big\{\mathrm{Per}(\Omega):\Omega\subset M,{\rm Vol}_{g}(\Omega)=t\,{\rm Vol}_{g}(M)\big\},$
@@
Each gap theorem above strengthens well-known rigidity theorems for spherical suspensions. For (1) use the maximal-diameter theorem of Cheng [Che75], for (2) use the Lichnerowicz-Obata theorem [Lic58, Oba62], for (4) use the Lévy-Gromov estimate [Gro81] (see a survey [Ant25]), for (5) use Bishop-Gromov based at the singular point [BC64, Gro81]. In the RCD setting, equality yields a spherical suspension; see for instance [Ket15, CaM17]. Note also that the $L^{1}$ smallness in (3) is implied by the $L^{p}$ smallness with any $p\geqslant 1$ and that gap theorems for smooth Einstein $n$-manifolds with small $L^{\frac{n}{2}}$ norm of their Weyl tensors are proven in [Sin92, She90, IS02].
@@
Direct generalizations of Corollary 1.3 fail in dimension $5$ and higher: in [Böh98], Böhm constructed a sequence of Einstein $5$-metrics $(S^{5},g_{i})$ with ${\rm Ric}(g_{i})=4g_{i}$ and volume uniformly bounded below such that $(S^{5},g_{i})_{i\in\mathbb{N}}$ converges in the Gromov-Hausdorff sense to a spherical suspension where the diameter and the first eigenvalue of the $(S^{5},g_{i})$ are arbitrarily close to $\pi$ and $5$ respectively, and supposedly, since the singularities forming along the sequence are conical, we should have: for $1\leqslant p<\frac{5}{2}$, $\|W\|_{L^{p}}\to 0$. See Question 1.31 for a reasonable higher-dimensional generalization.
@@
The constants in our gap theorems of Corollary 1.3 are not explicit since they are proven by contradiction. Gap theorems with explicit constant, such as [Gur00] are exceptionally rare; see [AES19].
@@ 1.2.3. Singular set of noncollapsed limits of Einstein $5$-manifolds, and orbifold structure
where $\mathcal{S}_{\Gamma}^{1}$ is the set of points whose unique tangent cone is $\mathds{R}\times\mathds{R}^{4}/\Gamma$, $\Gamma\subset O(4)$ and $\mathcal{S}^{0}$ is a set of countably many points where no tangent cone splits.
Our first result shows that locally each $\mathcal{S}^{1}_{\Gamma}$ is contained in the bi-Lipschitz image of some subset of $\mathds{R}$. In particular, this refines the classical rectifiability statement of [CJN21, JN21], eliminating the need to discard an $\mathcal{H}^{1}$-measure zero subset:
@@
Let $(M_{i},g_{i},p_{i})_{i\in\mathbb{N}}$ be a sequence of pointed $5$-manifolds with $|{\rm Ric}|<4$ and ${\rm Vol}(B_{1}(p_{i}))>v>0$. Let $(X,d,p)$ be a Gromov-Hausdorff limit of the sequence. For any $x\in\mathcal{S}^{1}_{\Gamma}$, there exists some $r>0$ and some function $\phi$ such that $\phi:B_{r}(x)\cap\mathcal{S}^{1}_{\Gamma}\to\mathds{R}$ is bi-Lipschitz. In particular, the singular set $\mathcal{S}$ satisfies: There exist countably many points $q_{j}$ and measurable subsets $Z_{i}\subset\mathcal{S}^{1}_{\Gamma_{i}}$ with bi-Lipschitz maps $\phi_{i}:Z_{i}\to\mathds{R}$ such that
@@
Corollary 1.19.
Each connected component $\Omega$ of $\mathcal{S}^{1}_{\Gamma}$ is either a single point or a geodesic, and thus locally convex. That is, for every $x\in\Omega$ there exists $r>0$ such that for any $y\in\Omega\cap B_{r}(x)$, every minimizing geodesic $\gamma$ joining $x$ and $y$ is contained in $\Omega$.
Remark 1.20.
In general, the subset $\mathcal{I}\subset I$ in Theorem 1.18 is non-empty and one cannot expect convexity of the entire singular set $\mathcal{S}$. For example, let $C(Z)$ be a Ricci-flat cone that does not split, where $Z$ is an Einstein $4$-orbifold with at least two singular points. Then the singular set is not locally convex around the vertex.
Theorem 1.18 characterizes any curve of singularities up to a nowhere dense subset: they are open geodesic segments. We now turn to the converse and completely characterize neighborhoods of open geodesics of our limit spaces.
@@
Let $\sigma:(-\varepsilon,\varepsilon)\to X$ be a geodesic in a noncollapsed Einstein $5$-dimensional limit space $(X,d)$ with $|{\rm Ric}_{M_{i}}|\leqslant 4$. Then either $\sigma\big((-\varepsilon,\varepsilon)\big)$ is regular, or the entire $\sigma\big((-\varepsilon,\varepsilon)\big)\subset\mathcal{S}^{1}_{\Gamma}$ for some $\Gamma\subset O(4)$ and the interior of the curve is removable in the sense of Theorem 1.18.
@@
Finally we discuss the orbifold regularity of the limit space $X$. Theorem 1.18 says the “interior” of each $\mathcal{S}^{1}_{\Gamma}$ is locally a geodesic along which the metric extends real-analytically. Then it is natural to introduce the orbifold-regular-singular decomposition as follows:
@@
In particular, we show that the orbifold-regular set $\mathcal{R}_{\mathrm{orb}}$ is homeomorphic to a real-analytic variety, thus confirming a conjecture of Naber [Nab14, Conjecture 6.4] away from extreme points in dimension $5$.
1.2.4. Singular set of higher dimensional noncollapsed limits
In higher dimensions we can distinguish the (n4)(n-4)-stratum from the (n5)(n-5)-stratum, and this separation already forces a uniqueness phenomenon.